A silversmith has two alloys, one containing 60% silver and the other 40% silver. How much of each should be melted and combined to obtain 800 grams of an alloy containing 54% silver?
Answer by mananth(16949)

Silver percent ---------------- quantity
Metal alloy I 60 ---------------- x grams
Metal alloy II 40 ------ 800 - x grams
Mixture 54 ---------------- 800

60x+ 40(800 -x)=54 * 800
6000x +4000(800-x) = 4320000
6000x+ 3200000 -4000x = 4320000
6000x- 4000x= 4320000 - -3200000
2000 x = 1120000
/ 2000
x = 560 grams 60.00% Metal alloy I
240 grams 40.00% Metal alloy II

m.ananth@hotmail.ca